OpenStudy (anonymous):

simplify 5^-8 * 5^10

jimthompson5910 (jim_thompson5910):

5^(-8)*5^10 = 5^(-8+10) = 5^2 = 25

OpenStudy (anonymous):

Its 25

OpenStudy (anonymous):

When multiplying exponents with like bases, you really are adding the exponents to get new exponent. 5^(-8+10) equals 5^2 or 25.